OpenStudy (anonymous):

Which of the following is a factor of 32x3 + 4? 2x – 1 4x2 – 4x + 1 4x2 – 2x + 1 None of the above

OpenStudy (anonymous):

4x2-2x+1

OpenStudy (amistre64):

if the value of x that makes the option go zero, makes the equation itself go zero; its a factor

OpenStudy (anonymous):

Thanks (:

OpenStudy (anonymous):

welcome

OpenStudy (amistre64):

2x – 1 = 0 when x=1/2 if the equation goes zero for x=1/2 its a factor 32x3 + 4 1/2 ---------- 32/8 + 4 = 4+4 = 8 .... first one aint a factor